# Diffusioosmotic corner flows

### v2

@@ VII Solution of the biharmonic equation for given solute concentration field
Figure 8: Diffusioosmotic corner flow for the wedge opening angle of $\pi/2$. (a) Isolines of the stream function show a single vortical structure in the corner. The active patches covers both walls close to the tip (for $\rho<1$), with emission at the horizontal wall and absorption at the vertical surface. (b) Transversal velocity profile along the bisector line of $\theta=\pi/4$.
@@ VIII Diffusion with analytic Neumann boundary conditions
$c=\lambda+\sum_{n\geq 1}(F_{n}\cos(n\theta)+G_{n}\sin(n\theta))\rho^{n},$ (63)
@@
for $n\geq 1$. While $\lambda$ remains arbitrary, it can be fixed by an additional boundary condition, such as the value of $c(\rho=0)$. Then, we get $c(\rho=0)=\lambda$. The condition for existence of $F_{n}$ and $G_{n}$ such that boundary conditions can be satisfied by this ansatz is $\sin(n\alpha)\neq 0$ (i.e. $\alpha\neq q\pi$, where $q\in\mathbb{Q}^{+}$). If $\sin(n\alpha)=0$, a more general ansatz is required. Alternatively, we note that the above ansatz is sufficient if $\alpha=r\pi$ for $r\in\mathbb{R}^{+}\backslash\mathbb{Q}^{+}$. Then $r$ can taken to be, for example, $r=q\pm\sqrt{2}\cdot 10^{-n}$, where $n$ can be arbitrarily large; Then we obtain a solution for the desired angle $\alpha=q\pi$ in the limit of large $n$.
@@ IX Discussion and conclusions
In this paper, inspired by Moffatt eddies [49, 50] that emerge in wedge-like geometries under mechanical boundary forcing, we have shown that corner geometry with phoretically active sectors can be used to generate similar eddies. Our solution involved the exact solution of the diffusion problem in the wedge geometry using the Mellin transform formalism, followed by a solution of the biharmonic equation for the stream function of the associated flow. The proposed diffusioosmotic framework relies on a fixed-flux boundary condition for the solute release from a catalytic patch, which is appropriate when the diffusion of the solute is much faster than its adsorption (that is, in the limit of low Damköhler number) and that the advection is slow compared to diffusion (corresponding to the limit of low Péclet number).
In the simple case of uniform coverage of the walls with chemical activity, the geometric asymmetry alone is responsible for heterogeneous distribution of solute concentration, which drives the flow. In such a setting, the problem admits a simple analytical solution akin to the classical Taylor’s scraper problem when one wall is chemically active. Similarly, an analytical solution is also available when both walls are covered with catalyst. These solutions can serve as benchmarks for numerical computations, particularly when solving for the concentration field with non-trivial phoretic boundary conditions.
In the more general case, where patches of activity are distributed on the walls, we presented a solution technique using the Mellin transform to solve the diffusion problem and, subsequently, find the coefficients of the general solution of the associated flow problem. The solutions can be evaluated numerically for an arbitrary distribution of patches and combination of boundary conditions.
@@
Data availability statement The Mathematica notebooks used to compute the results of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.16938258, reference number 16938258.

### v3

@@ VII Solution of the biharmonic equation for given solute concentration field
Figure 8: Diffusioosmotic corner flow for the wedge opening angle of $\pi/2$. (a) Isolines of the stream function show a single vortical structure in the corner. The active patches cover both walls close to the tip (for $\rho<1$), with emission at the horizontal wall and absorption at the vertical surface. (b) Transversal velocity profile along the bisector line of $\theta=\pi/4$.
@@ VIII Diffusion with analytic Neumann boundary conditions
$c={\lambda_{0}}+\sum_{n\geq 1}(F_{n}\cos(n\theta)+G_{n}\sin(n\theta))\rho^{n},$ (64)
@@
for $n\geq 1$. While $\lambda_{0}$ remains arbitrary, it can be fixed by an additional boundary condition, such as the value of $c(\rho=0)$. Then, we get $c(\rho=0)={\lambda_{0}}$. The condition for existence of $F_{n}$ and $G_{n}$ such that boundary conditions can be satisfied by this ansatz is $\sin(n\alpha)\neq 0$ (i.e. $\alpha\neq q\pi$, where $q\in\mathbb{Q}^{+}$). If $\sin(n\alpha)=0$, a more general ansatz is required. Alternatively, we note that the above ansatz is sufficient if $\alpha=r\pi$ for $r\in\mathbb{R}^{+}\backslash\mathbb{Q}^{+}$. Then $r$ can taken to be, for example, $r=q\pm\sqrt{2}\cdot 10^{-n}$, where $n$ can be arbitrarily large; Then we obtain a solution for the desired angle $\alpha=q\pi$ in the limit of large $n$.
@@ IX Discussion and conclusions
In this paper, inspired by Moffatt eddies [47, 48] that emerge in wedge-like geometries under mechanical boundary forcing, we have shown that corner geometry with chemically active sectors can be used to generate similar eddies. Our solution involved the exact solution of the diffusion problem in the wedge geometry using the Mellin transform formalism, followed by a solution of the biharmonic equation for the stream function of the associated flow. The proposed diffusioosmotic framework relies on a fixed-flux boundary condition for the solute release from a catalytic patch, which is appropriate when the diffusion of the solute is much faster than its adsorption (that is, in the limit of low Damköhler number) and that the advection is slow compared to diffusion (corresponding to the limit of low Péclet number).
In the simple case of uniform coverage of the walls with chemical activity, the geometric asymmetry alone is responsible for heterogeneous distribution of solute concentration, which drives the flow. In such a setting, the problem admits a simple analytical solution akin to the classical Taylor’s scraper problem when one wall is chemically active. Similarly, an analytical solution is also available when both walls are covered with catalyst. These solutions can serve as benchmarks for numerical computations, particularly when solving for the concentration field with non-trivial active boundary conditions.
In the more general case, where patches of activity are distributed on the walls, we presented a solution technique using the Mellin transform to solve the diffusion problem and, subsequently, find the coefficients of the general solution of the associated flow problem. The solutions can be evaluated numerically for an arbitrary distribution of patches and a combination of boundary conditions.
We note here that while our present work bears resemblance to the previously analysed case of flow in a 2D confined channel of Visan et al. [77] in the limit of $\alpha\to\pi$ with a single active sector, a straightforward comparison is not possible. In our case, there is a single wall, and the concentration field decays monotonically with the distance from the active patch, while Visan et al. [77] consider a system confined between two walls; the bottom surface is patterned with activity, and they impose a Dirichlet boundary condition of constant concentration on the top wall. Additionally, they assume periodic boundary conditions in the direction parallel to the walls. They then observe a pair of counter-rotating vortices. In our system, with only one surface being present for $\alpha=\pi$, we see similar flow structure in the vicinity of an active patch. However, the presence of the top boundary in their system introduces an additional length scale (apart from the size of the active patch itself) that sets the size of recirculating vortices. This interplay of length scales is further discussed by Michelin & Lauga [43]. In the absence of the top confining wall, we see different asymptotics of the concentration field (decay with the distance), and the recirculation zone becomes larger but the flow strength decays faster away from the patch.
@@
We note that our solution can also be applied as an approximation in finite geometries, such as that of a rectangular microfluidic channel with chemically patterned corners. When active sectors are contained in the region of a corner (with $\alpha=\pi/2$) such that $\rho<b_{\text{max}}$, the concentration field away from the active region asymptotically behaves as $c(\rho,\theta)\sim\left({b_{\text{max}}}/{\rho}\right)^{2}$. Thus, the slip velocity due to chemically active sectors scales as $|\bm{u}_{s}|\sim\left({b_{\text{max}}}/{\rho}\right)^{3}$. Note that for such approximation to apply to finite domains, total solute flux away from the chemically active region has to vanish. It can be verified that eq. (60) is a sufficient condition for the total flux to vanish for $\rho>b_{\text{max}}$, and the local solute current asymptotically scales as $\left({b_{\text{max}}}/{\rho}\right)^{3}$. In the limit of well-separated chemically active corners, one can superpose the solutions from neighbouring corners, and as a result produce the streamline pattern for a more complex channel shape.
@@
Data availability statement The Mathematica notebooks used to compute the results of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.17800756, reference number 17800756.